# The Three Regimes of Offline-to-Online Reinforcement Learning

### v2

@@ C.1 Q-Function-Based Taxonomy
Conservative offline RL aims to outperform the behavior policy that generates the offline dataset (levine2020offline). Ideally, the value function of the pretrained policy $\pi_{0}$ should therefore exceed that of the behavior policy $\pi_{\mathcal{D}}$ on in-dataset state-action pairs. Formally, kostrikov2021offline show that for in-sample Q-learning, the optimal value function satisfies
@@
Table 14: Confusion matrix of fine-tuning results using Q-based taxonomy. Green cells: correct predictions (32/63); red cells: opposite mismatches (19/63); gray cells: adjacent mismatches (12/63). Overall, Q-based taxonomy achieve 51% correct predictions with 30% opposite mismatches.
@@
Fine-tune
$\pi_{0}$-centric >> $\mathcal{D}$-centric 23 0 4
@@
C.2 Behavior-Cloning-Based Taxonomy
@@
Fine-tune
$\pi_{0}$-centric >> $\mathcal{D}$-centric 17 10 0

### v3

@@ C.2 Q-Function-Based Taxonomy
Conservative offline RL aims to outperform the behavior policy that generates the offline dataset (Levine et al., 2020). Ideally, the value function of the pretrained policy $\pi_{0}$ should therefore exceed that of the behavior policy $\pi_{\mathcal{D}}$ on in-dataset state-action pairs. Formally, Kostrikov et al. (2021, Lemma 2) show that for in-sample Q-learning, the optimal value function satisfies
@@
Table 17: Confusion matrix of fine-tuning results using Q-based taxonomy. Green cells: correct predictions (32/63); red cells: opposite mismatches (19/63); gray cells: adjacent mismatches (12/63). Overall, Q-based taxonomy achieves 51% correct predictions with 30% opposite mismatches.
@@
Fine-tune $\pi_{0}$-centric >> $\mathcal{D}$-centric 23 0 4
@@
C.3 Behavior-Cloning-Based Taxonomy
@@
Fine-tune $\pi_{0}$-centric >> $\mathcal{D}$-centric 17 10 0